One number is six less than a second number. Six times the first is 3 more than 3 times the second. Find the numbers
Answer by solve_for_x(190)

Let x represent the first number, and x + 6 represent the second number.

Then "six times the first is 3 more than 3 times the second" can be written as the
following equation:

6x = 3(x + 6) + 3

Expanding the terms on the right side gives:

6x = 3x + 18 + 3

6x = 3x + 21

6x - 3x = 21

3x = 21

x = 7

The first number is 7.

The second number is x + 6 = 7 + 6 = 13.